OpenStudy (anonymous):

x^2+y^2+8y=0 find the radius and the center and graph

OpenStudy (anonymous):

the same procedure in your last question. complete the square of the variable y. follow the general formula for a circle. you should be able to get the center and the radius.

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

okay. I'll show it step by step. as i said, the general formula is (x - h)^2 + (y - k)^2 = r^2 the equation is this, x^2 + y^2 + 8y = 0 you need to complete the square of the variable y. so it would look like this, x^2 + y^2 + 8y + 16 = 16 then, transform the equation, (x+0)^2 + (y+4)^2 = 16 you can find now the center and the radius. from the general formula, the center is (h,k) and the radius is r